In the least-squares line
ŷ = 5 − 8x,
what is the value of the slope?


When x changes by 1 unit, by how much does ŷ change?
a.)When x decreases by 1 unit, ŷ decreases by 8 units.
b.)When x increases by 1 unit, ŷ decreases by −8 units.
c.)When x increases by 1 unit, ŷ decreases by 8 units.
d.)When x increases by 1 unit, ŷ increases by 8 units.

In the least-squares line ŷ = 5 + 6x, what is the marginal change in ŷ for each unit change in x?
1
Expert's answer
2019-11-18T12:27:57-0500

Given "\u0177 = 5 \u2212 8x," the slope is given by "\frac{dŷ}{dx}=\frac{d(5-8x)}{dx}=\frac{d(5)}{dx}−8⋅\frac{dx}{dx}=0−8=−8". When x increases by 1 unit, ŷ decreases by 8 units. Equivalently, when x decreases by 1 unit, ŷ decreases by 8 units.

Given "\u0177 = 5 + 6x", marginal change in ŷ for each unit change in x is given by "\frac{dŷ}{dx}=\frac{d(6x+6)}{dx}=6\frac{dx}{dx}+\frac{d(5)}{dx}=6+0=6". Thus, a unit change in x leads to 6 units change in ŷ. That is a unit increase in x leads to 6 units increase in ŷ.